A - 1.50 m A/3 A/3 B/3 8/3 -steal gael - S 355 - overall olepth of slab 155 mm Load due to floor fnishes o0 2o KNIm -lmPosed Load: 3.00 kNlmt Flosr slab overall ddepth of slab 160 mm Load oue to flost inishes 10 NI Complete the following structural analysis and design tasks complying with the BS EN: 1993-1 code specifications.
Assume that the compression flange of the Plate Girder is fully restrained. The Plate Girder is supported on a heavy stiffened bracket at each end.
i. Determine dimensions of the plate girder and section classification
Answer #1

Step 1 Calculation of Factored load

for roof slab:

total load of roof slab = load due to roof finish + imposed load

= 3.0+0.8

= 3.8 kN/m2  

now Factored load (due to roof slab ) = 3.8 x 1.5

= 5.7 kN/m2

for floor slab

total load of floor slab = load due to floor finish + imposed load

= 1.0+1.75

= 2.75 kN/m2  

now Factored load (due to floor slab ) = 2.75 x 1.5

= 4.125 kN/m2

Total Factored load (due to roof and floor)= 5.7 + 4.125

= 9.85 kN/m2

Step 2: Calculation of maximum Bending Moment(BM) and Shear force(SF)

Maximum BM = WL2/8 = 9.85x152/8

= 277 kN.m

(as plate girder will be placed along theloanger span so l = 15)

Maximun SF = WL/2 = 9.85x15/2

= 73.9 kN

STEP 3: Depth of Web portion

Dw = 5 ( BM/fy)1/3

= 5 x (277/355)1/3 ( fy= Grade of steel)

Dw = 4603mm = 4.6 m

STEP 3: Thickness of web

for stiffened brackets condition below should satisfy

Dw/tw<100

Take Dw/tw< 90   

tw = 4603/ 90

= 5114mm = 5.2 m

STEP 5 Area of flange

Af= 15 x 7.5

= 112.5 m2

STEP 6 Width of Flange (bf)

bf=0.25 dwto 0.3 dw

bf= 0.25x4603 to .30x4603

bf= 1150 to 1380

provide bf= 1200 mm

STEP 7 Thickenss of flange

tf=Af/bf

tf= 112.5/12

tf= 9.375 m

STEP 8 Selection of section

According to the calculated data choose the section from the steel table

so providing plates for web and flange and providing welded connection at preffered spacing

(let the weldiing is done at 50mm spacing with 50mm weld)

choosing plate or flats and welding the together will give plate girder